a wood block of mass m=300g rests on a larger wood block of mass m=4kg that rests a wooden table.the coefficients of static and kinetic friction between al surfaces are 0.4 and 0,2 respectively.a horizontal force ,f, is applied to the lower block
1
Expert's answer
2020-12-04T11:16:13-0500

Answer

Mass of upper block(m) =0.3kg

Mass of lower block(M) =4kg

If horizontal F force applied on lower block then it's minimum value should be

"F=\mu _1(m+M) g+\mu_2mg\\F=0.2(0.3+4) \times 9.81+0.4\times 0.3\times9.81\\F=9.61N"